This website was created as part of Berkeley Lab's emergency response plan for local and regional disasters. In the event of any disruption of services, this site will provide a means for Lab employees to communicate with each other and stay informed on the status of lab operations. Internet requests for will automatically redirect to this site if an emergency plan has been put into effect. 

The site can also be used by customers and lab employees who are having difficulty connecting to lab services to verify that no disruption has taken place and all services are available. The Emergency Status Information Line is: 1-800-445-5830

The Laboratory is under NORMAL operations and all services are available.


11 October 2017, 8:06 AM : Laboratory LDAP Sign-in disrupted for some users--RESOLVED

7:40 AM : Laboratory LDAP Sign-in disrupted for some users

The IT Division is responding to user reports of problems logging into systems using LDAP authentication.  This problem was first reported early this morning and we will provide updates as we diagnose the problem.


October 4, 2017 5AM - A power issue is impacting 50a, 50c, and 50e and these buildings are currently without power.   Facilities is investigating.    Buildings 50 and 50B currently have power, but did experience a voltage sag during the event. Access to these buildings is currently restricted (7AM)

Thursday, September 7 Disruption with Google Drive Service for some users -- RESOLVED by Google 0924 AM

As of 0737 AM PT, Google is reporting an issue with Google Drive. 
Google Drive is not loading files and results in a 500 error for some users.  The IT-Division will update this site as we learn more or the problem is resolved by Google.

Ongoing: Power Outage and IT Outage Aug 18-21 2017 -- Resolved


All IT services are restored and available.

Scientific Computing is working with staff impacted by loss of data on the global/scratch file system.  

No further updates to follow on for Aug 18-21 Power Outage.


All IT services are restored and available except for Lawrencium's global/scratch file system. Restoration of global/scratch is expected for Thursday AM, and users will be notified via email when this occurs.


Most of the Scientific Computing cluster infrastructure is back online but Lawrencium's global/scratch file-system remains offline. The issue has been escalated with the vendor and the team will provide an update tomorrow end of day.  

Other IT services are restored and available.


Most of the Scientific Computing cluster infrastructure is back online but Lawrencium's global/scratch file-system is still offline. The issue has been escalated with the vendor and the team will provide an update late tonight or in the morning.  

Other IT services are restored and available.


Most services restored, however Scientific Computing Lawrencium scratch-storage availability delayed.  The team is currently working on this problem with the vendor and will update status later this afternoon.  Current best estimate for return to service of this storage is end of day August 21.


SW A6 Electrical Maintenance Phase 1 - Saturday through Sunday - August 19th & 20th, 2017 

We will turn off the power on Saturday around 5am and restore power on Sunday around 10pm

What is to be accomplished: 3 year High Voltage Electrical Distribution Preventative Maintenance

What will be impacted: Building Electrical Power (similar to a PG&E power outage)

What will not be impacted: Standby Emergency Generator Power and all equipment connected to it.

(i.e. Critical exhaust systems for Fume Hoods, Critical Freezers, Life Safety Systems, etc.)


Buildings that will lose normal power:

50 Complex, 54 (Café), 65 Complex, 70, 70A, 88, and the Blackberry Gate 


Prior to the outageWe recommend that everyone turn off his or her computers and other sensitive equipment.

During the outage Buildings will be dark.  The HVAC systems and card key access systems will not be working (fail in locked position).  You will not be allowed to enter the affected buildings unless you are an essential participant in the maintenance project and have received permission from the Project Manager.


Important Information Notes:


1. All equipment currently connected to standby emergency generator power will remain energized during the shutdown.



Maintenance to LBL power systems impacting Building 50 will take place over the weekend of Aug 19-20.

The following systems are expected to be operational:

  • Login / LDAP / Directory Services
  • Email delivery, routing, and access including email lists.
  • All Google Services
  • Building management and security systems operated for Facilities
  • Telephones and Voicemail
  • Network access to unaffected buildings on the LBL site.


The following systems will be taken offline:

  • All IT Operated Business Systems (Peoplesoft Finance, Peoplesoft HR, LETS, Maximo, etc).
  • Commons (Confluence) (will include websites for facilities, IT and other sites hosted on this platform).
  • including all sites hosted on this platform.
  • All HPC Services
  • All IT Provided CIFS (Home Directory, Group Share) Services.
  • All services in 2109 Colocation Facility.


Systems will have different restoration points, but restoration could easily extend into Monday August 21st.

Please monitor for updates.

Previous Events

August 2-3:  Fire and Power Event.  Services restored.

May 9-10 2017: Major sitewide power outage.  Services restored.

March 13 2017:  Unscheduled outage of outbound and inbound calling.  Service restored.

Feb 21 2017:  7PM Multiple server outages for scheduled maintenance including commons.   Expected return to service by 10PM tonight.

May 8, 2016: Maintenance for Financial System May 6 Noon - May 8 8AM 2016

April 25, 2016:  Voicemail system outage.  Lab users are unable to leave or retrieve Lab voicemails - RESOLVED

Feb 18, 2016: Unscheduled major network outage impacting many parts of LBL.  - RESOLVED

Dec 18, 2015:  Post 5pm minor outages scheduled for, email lists, and other service.  More details here>>>. - RESOLVED

Dec 27-29, 2015:  Numerous small outages scheduled for offsite buildings and substantial outages scheduled on the main site for Dec 28th.  More details here>>>>  - RESOLVED

Dec 25, 2015: Unscheduled outage of identity management (login, ldap, AD, Pledge/OTP etc) - RESOLVED

Nov 7, 2015: Onsite Power Outage.  Details here>>>  - RESOLVED

Oct 15, 2015:,, problem resolved. 1:03pm.  - RESOLVED

Oct 15, 2015: We are experiencing an outage on,, and starting at approximately 12:15pm. The service provider is in the process of restoring those sites. We will update this page when new information is available.  - RESOLVED

Oct 9, 2015:  Drive problem resolved. 1:40pm  - RESOLVED

Oct 9, 2015: Google Drive is experiencing an outage which started at approximately 11:30am. Drive is coming back as of  1:15pm, but access to sheets continues to be an issue.  - RESOLVED

July 11, 2015 Power Outage  - RESOLVED

Feb 27, 2015 Power Outage  - RESOLVED

Feb 2, 2015: Scheduled outage for,, and other hosted wordpress sites - RESOLVED


Aug 24, 2014: Following this morning's 6.0 Napa earthquake, Berkeley Lab staff are conducting a coordinated damage assessment of the Lab site. --Resolved

July 11, 2014 2PM:  We're making changes to,, and - RESOLVED

May 16, 2014: Scheduled downtime for www and several other applications until 9pm. - RESOLVED

Mar 14, 2014  Minor Scheduled Maintenance 6-9pm, and selected additional systems.- RESOLVED

Jan 17, 2014  Major outage of IT Systems scheduled for Friday Jan 17th 6pm-10pm  More    Most services restored by 8.30PM- RESOLVED

Dec 30, 2013 at 12:15PM there was  a problem during a scheduled electrical maintenance activity.  Connectivity was restored at 4:05PM. - RESOLVED

Dec 13, 2013 at 6PM - Major outage of IT provided services.   information>>>- RESOLVED

Nov 6, 2013:  The power distribution hardware described in the incident this morning failed again around 12:30PM today, impacting www, mail delivery, commons, and many other services.  It was resolved around 12:48.- RESOLVED

Nov 6, 2013: We experienced an outage of www, commons, and email delivery due to the failure of a piece of power distribution hardware in the datacenter.  Services have been restored. - RESOLVED

April 2, 2012: LBNL experienced technical problem with the main Lab web server ( from 6:30pm to 7:00pm.- RESOLVED

July 26, 2012: We conducted a test of our website disaster recovery procedure from 5:00pm to 5:15pm Thursday, July 26, 2012. During the test, users attempting to access were redirected to this site.- RESOLVED

July 5, 2012: We conducted a test of our website disaster recovery procedure from 7:00pm to 7:15pm Thursday, July 5, 2012. During the test, users attempting to access were redirected to this site.- RESOLVED

June 8, 2012: Due to an emergency outage, Berkeley Lab experienced a loss of Internet connectivity. Email, voicemail, and other IT services were impacted.- RESOLVED

Dec 2-3 2011 Outage is Complete.  All systems are operational.  - RESOLVED

For Google Apps for Berkeley Lab Status Information Click Here.
 Note that this link shows status information for all Google Apps services, not just those provided to LBL, so an issue may display there that doesn't impact us.

Recent Networking Events Are Recorded Here: